# Boundary conditions matter: On the spectrum of infinite quantum graphs

### v2

@@ 2.4. Laplace-type operators on infinite graphs
$\operatorname{dom}(\mathcal{H})=\left\{f\in\operatorname{dom}(\mathcal{H}_{\max})\cap L^{2}(\mathcal{G}):f\in C(\mathcal{G})\text{ and }\sum_{\mathsf{e}\in\mathsf{E}_{\mathsf{v}}}\frac{\partial}{\partial\nu_{\mathsf{e}}}f(\mathsf{v})=0\text{ for all }\mathsf{v}\in\mathsf{V}\right\}$
@@
(2.6) $\operatorname{\mathfrak{t}}_{\mathrm{F}}[f]$ $:=\int_{\mathcal{G}}|f^{\prime}(x)|^{2}\,\mathrm{d}x,$ $\operatorname{dom}[t_{\mathrm{F}}]:=H^{1}_{0}(\mathcal{G}).$
@@
It can be shown, see [37, Corollary 6.7] and also Remark 2.16.(3), that on metric graphs with a finite number of ends of finite volume, $\mathcal{H}_{\mathrm{N}}$ corresponds to imposing Neumann boundary conditions at all ends while the Friedrichs realisation $\mathcal{H}_{\mathrm{F}}$ imposes Dirichlet boundary conditions at all ends.
@@ 3.1. Embedding theorems for Sobolev spaces
This section is about criteria for discreteness of the spectrum of the operator $\mathcal{H}_{\mathrm{N}}$ and all self-adjoint Laplacian realisations dominating $\mathcal{H}_{\mathrm{N}}$ in the sense of forms, including $\mathcal{H}_{\mathrm{F}}$. Since discreteness of the spectrum of self-adjoint realisations is equivalent to compactness of the embedding of their form domains in $L^{2}(\mathcal{G})$ this subsection can also be understood as discussing Sobolev embedding theorems. For (locally finite) infinite metric graphs with compact metric completion it has been known since [14, Theorem 2.2] that $H^{1}(\mathcal{G})$ embeds compactly into $C_{b}(\overline{\mathcal{G}})$. We stress that, however, this does not necessarily imply a compact embedding into $L^{2}(\mathcal{G})$, unless $\overline{\mathcal{G}}$ has finite volume (for a counterexample, see the diagonal comb graph in 3.4.(2)).
In particular, the resolvent of $\mathcal{H}_{\mathrm{N}}$ and hence – by pointwise domination of the integral kernels – also of $\mathcal{H}_{\mathrm{F}}$ or any Markovian realisation – will be trace class operators, see [37, Theorem 5.1]. However, this finite volume criterion is insufficient to characterise discreteness of the spectrum: there are metric graphs of infinite volume on which $\mathcal{H}_{\mathrm{F}}$ and even $\mathcal{H}_{\mathrm{N}}$ have purely discrete spectrum: Non-trivial examples are given in [2, Theorem 1.3.5] and [59, Theorem 4.1].
@@
The subgraph $\mathcal{G}_{c}^{*}$ can be identified with an Eulerian walk of finite length $p$ which allows to identify $*f^{*}_{|\mathcal{G}_{c}^{*}}\in B^{*}$ with a function $\hat{f}\in H^{1}([0,p])$. On $\mathcal{G}_{c}^{*}$ we can then define
@@
We cover $[0,p]$ with $N$ non-overlapping intervals of length $\delta$,
$[0,p]=\bigcup_{j=1}^{N}\overline{I_{j}},\quad\text{where}\quad I_{j}\cap I_{k}=\emptyset,\quad\text{for $k\neq j$},\quad\text{and}\quad\lvert I_{j}\rvert=\delta\quad\text{for all $j\in\{1,\dots,N\}$}.$
$P(f^{*})(x):=\begin{cases}\frac{1}{\delta}\int_{I_{j}}f^{*}(x)\mathrm{d}x\quad&\text{if $x\in\mathcal{G}_{c}^{*}$, and $x\in I_{j}$ on the Eulerian walk},\\ 0\quad&\text{if $x\in(\mathcal{G}_{c})^{\complement}$}.\end{cases}$
@@
and note that the first summand in (3.3) is at most $\varepsilon^{2}$ by assumption. The second term in (3.3) can be considered as an integral along the Eulerian walk and estimated by
$\sum_{i=1}^{N}\int_{I_{j}}\left\lvert\frac{1}{\delta}\int_{I_{j}}\left(f^{*}(x)-f^{*}(y)\right)\mathrm{d}y\right\rvert^{2}\mathrm{d}x\leq\sum_{i=1}^{N}\int_{I_{j}}\frac{1}{\delta}\int_{I_{j}}\left\lvert f^{*}(y)-f^{*}(x)\right\rvert^{2}\mathrm{d}y\ \mathrm{d}x.$
We substitute $h=y-x$ and emphasise that $h\in(-\delta,\delta)$. Thus, (3.2) implies
$\lVert f^{*}-Pf^{*}\rVert_{L^{2}(\mathcal{G^{*}})}^{2}$ $\leq\varepsilon^{2}+\sum_{i=1}^{N}\int_{I_{j}}\frac{1}{\delta}\int_{-\delta}^{\delta}\left\lvert f^{*}(x+h)-f^{*}(x)\right\rvert^{2}\mathrm{d}h\ \mathrm{d}x$
@@
Furthermore, $P$ is a bounded operator, and $B^{*}$ is bounded, thus $P(B^{*})$ itself is bounded. Finally, since $P$ is of finite rank, it is totally bounded. By [28, Lemma 1], we find that $B^{*}$ is totally bounded in $L^{2}(\mathcal{G}^{*})$, and thus the same is true of $B$ in $L^{2}(\mathcal{G})$.
@@
Next, we obtain a criterion on compactness of embeddings of Sobolev spaces on trees. It is a generalisation of results in [38], see for instance Lemma 8.1 therein, where infinite trees without degree one vertices are considered. We also mention that the metric condition in Proposition 3.2 is already known to play a role for spectral properties of infinite quantum graphs with $\delta$-couplings, see [21, Theorem 3.5].
@@
On infinite trees without degree one vertices (“leaves”), the statement follows immediately from Theorem 3.1. It remains to prove the statement in the presence of leaves. Compactness of the embedding is equivalent to
@@ 3.2. Spectral phase transition on graphs of infinite volume
For $\alpha>0$, we define the “diagonal comb” metric graph $\mathcal{G}_{\alpha}$ by taking the interval $(0,1]$ (“horizontal shaft”), putting on every point $\frac{1}{n^{\alpha}}$, $n\in\mathbb{N}$, a vertex, and attaching to it an edge (“tooth”) of length $\frac{1}{n^{\alpha}}$.
001111
@@
Figure 3.1 contains an illustration in the case $\alpha=\frac{1}{2}$. Note that the larger $\alpha$, the sparser the teeth become. In particular, if $\alpha>1$, the graph has finite volume in which case $\mathcal{H}_{\mathrm{N}}$ (or any Markovian realisation of $\mathcal{H}$) will have purely discrete spectrum by [14, Theorem 2.2].
@@
    If $\alpha\in(0,\frac{1}{2}]$, then there is an $L^{2}(\mathcal{G}_{\alpha})$-orthonormal sequence of $H^{1}(\mathcal{G}_{\alpha})$-functions with uniformly bounded $H^{1}(\mathcal{G}_{\alpha})$ norm. In particular, $\mathcal{H}_{\mathrm{N}}$ has nonempty essential spectrum.
@@
Note that the phase transition happens at $\alpha=1/2$, that is, among graphs of infinite volume (the transition from infinite to finite volume being at $\alpha=1$); see also Remark 3.6. Also, all $\mathcal{G}_{\alpha}$ are trees with only a finite number of edges of length larger than $\varepsilon$ for any $\varepsilon>0$ – the criterion for discreteness of the spectrum of the Friedrichs realisation in Proposition 3.2. Thus, Theorem 3.4 also demonstrates that this criterion for discreteness of the Friedrichs realisation on trees does not extend to the Kirchhoff realisation.
Remark 3.5.
It is known that on any such domain, the spectrum of the Dirichlet Laplacian is purely discrete [55, 8]. Theorem 3.4.(3) or more generally the compactness criterion of Proposition 3.2 seem to be the corresponding analogues on metric graphs. However, for the Neumann Laplacian on domains, the situation is different. Indeed, on connected domains of infinite volume, the essential spectrum of the Neumann Laplacian is always nonempty, see the appendix of [18], which also provides examples of horn-shaped domains of finite volume with nonempty spectrum. Thus, Theorem 3.4.(1)–(2), which describes a phase transition from purely discrete to nonempty essential spectrum for the Neumann realisation among metric graphs of infinite volume, seems to present a new, metric-graph specific phenomenon which has no obvious equivalent among domains.
@@
Denote by $\mathsf{e}_{k}$ the $k$-th tooth (counted from $1$ to 00, i.e. from right to left in the orientation of Figure 3.1), and note that all points in $\mathsf{e}_{k}$ are at most at distance $2k^{-\alpha}$ to 00; hence
@@
$\begin{cases}&\text{linearly rising from $0$ to $1$ in $[\frac{1}{\sqrt{2n}},\frac{1}{\sqrt{n}}]$ on the horizontal shaft $\mathsf{e}$,}\\ &\text{linearly falling from $1$ to $0$ in $[\frac{1}{\sqrt{n}},\frac{2}{\sqrt{n}}-\frac{1}{\sqrt{2n}}]$ on the horizontal shaft,}\\ &\text{constant on the teeth and and zero everywhere else.}\end{cases}$
@@
Since, after passing to a subsequence, we can make the $\phi_{n}$ have mutually disjoint supports, up to renormalisation they yield an $L^{2}(\mathcal{G})$-orthonormal sequence of $H^{1}(\mathcal{G})$ functions with uniformly bounded $H^{1}(\mathcal{G})$-norm. (3) This is a direct consequence of Proposition 3.2. ∎
@@
Theorem 3.4.(2) asserts that the essential spectrum is nonempty but provides no further information on its structure. However, a closer look at the proof shows that, for $\alpha$ strictly below the critical threshold $\alpha=\frac{1}{2}$, the infimum of the spectrum is zero, since the test functions constructed in the proof will then have Rayleigh quotients converging to zero. It would be interesting to understand $\inf\sigma(\mathcal{H}_{\mathrm{N}})$ in the critical case $\alpha=\frac{1}{2}$.
@@ 3.3. Spectral approximation
Since $\mathcal{G}$ is connected, and the semigroups generated by both $\mathcal{H}_{\mathrm{F}}$ and $\mathcal{H}_{\mathrm{N}}$ are positive and irreducible, standard Perron–Frobenius theory, see e.g. [4, Proposition 4.12], implies that $\mu_{1}$ and $\lambda_{1}$ are necessarily simple. If $\mathcal{G}$ has finite volume, then $\mu_{1}=0$ with corresponding eigenfunctions being the constant functions, while $\lambda_{1}>0$, as we shall see below. In order to emphasise the dependence on $\mathcal{G}$, we will also write $\mu_{k}(\mathcal{G})$ and $\lambda_{k}(\mathcal{G})$. In the latter case, if there is any danger of ambiguity we will also include the Dirichlet vertex set, $\lambda_{k}(\mathcal{G})=\lambda_{k}(\mathcal{G},\mathfrak{V})$.
@@
On the other hand, given any $u\in H^{1}_{0}(\mathcal{G})$, by (3.9) and the definition of $H^{1}_{0}(\mathcal{G})$ as the closure in $H^{1}$ of $H^{1}_{c}(\mathcal{G})$, there exists a sequence of functions $u_{n}\in H^{1}_{0}(\mathcal{G}_{n})$ such that $u_{n}\to u$ in $H^{1}_{0}(\mathcal{G})$. Fix $k\geq 1$, write $\varphi_{j}:=\varphi_{j}(\mathcal{G})\in H^{1}_{0}(\mathcal{G})$ for the normalised eigenfunctions for $\lambda_{j}(\mathcal{G})$, $j=1,\ldots,k$, which will be our $u$ and for each $j$ choose $u_{j,n}\in H^{1}_{0}(\mathcal{G}_{n})$ such that $u_{j,n}\to\varphi_{j}$ in $H^{1}_{0}(\mathcal{G})$ (and hence also in $C(\overline{\mathcal{G}})$). Note that, by the dominated convergence theorem (using, e.g., $2|\varphi_{i}\varphi_{j}|\in L^{\infty}(\mathcal{G})$ as a dominating function, cf. [37, Lemma 3.2]), for all $i\neq j$,
@@
By induction on the trees $\mathcal{T}_{1},\ldots,\mathcal{T}_{j}$, it suffices to show that $\mu_{k}(\mathcal{G})\leq\mu_{k}(\mathcal{G}\setminus\mathcal{T}_{1})$ for all $k\in\mathbb{N}$. But this, in turn, follows from an argument completely analogous to the proof of [10, Theorem 3.10(1)] (with $r=1$): we first observe that clearly
@@
for all $k\in\mathbb{N}$, since the spectrum of the disjoint union of $\mathcal{T}_{1}$ and $\mathcal{G}\setminus\mathcal{T}_{1}$ is equal to the union of their spectra (counting multiplicities), and $\mu_{1}(\mathcal{T}_{1})=0\leq\mu_{k}(\mathcal{G}\setminus\mathcal{T}_{1})$ for all $k\in\mathbb{N}$. Next observe that $\mathcal{G}$ is formed from $\mathcal{T}_{1}\dot{\cup}(\mathcal{G}\setminus\mathcal{T}_{1})$ by gluing the two graphs together at the vertex $\mathsf{v}_{1}$; at the level of $H^{1}$-spaces, we have that $H^{1}(\mathcal{G})$ may be identified with the codimension one subspace of $H^{1}(\mathcal{T}_{1}\dot{\cup}(\mathcal{G}\setminus\mathcal{T}_{1}))$ consisting of those functions whose values satisfy an additional continuity condition at $\mathsf{v}_{1}$ (or rather its preimage vertices in $\mathcal{T}_{1}$ and $\mathcal{G}\setminus\mathcal{T}_{1}$ before gluing). The proof of [10, Theorem 3.4] may be repeated verbatim to give

### v3

@@ 2.4. Laplace-type operators on infinite graphs
$\operatorname{dom}(\mathcal{H})=\left\{f\in\operatorname{dom}(\mathcal{H}_{\max})\cap L^{2}(\mathcal{G}):f\in C(\mathcal{G})\text{ and }\sum_{\mathsf{e}\in\mathsf{E}_{\mathsf{v}}}\frac{\partial}{\partial\nu_{\mathsf{e}}}f(\mathsf{v})=0\text{ for all }\mathsf{v}\in\mathsf{V}\right\}.$
@@
(2.10) $\operatorname{\mathfrak{t}}_{\mathrm{F}}[f]$ $:=\int_{\mathcal{G}}|f^{\prime}(x)|^{2}\,\mathrm{d}x,$ $\operatorname{dom}[t_{\mathrm{F}}]:=H^{1}_{0}(\mathcal{G}),$
from which the Kirchhoff-type conditions at the vertices arise naturally.
@@
It can be shown, see [43, Corollary 6.7] and also Remark 2.17.(3), that on metric graphs with a finite number of ends, each of which of finite volume, $\mathcal{H}_{\mathrm{N}}$ corresponds to imposing Neumann boundary conditions at all ends while the Friedrichs realisation $\mathcal{H}_{\mathrm{F}}$ imposes Dirichlet boundary conditions at all ends.
@@ 3.1. Embedding theorems for Sobolev spaces
This section is about criteria for discreteness of the spectrum of the operator $\mathcal{H}_{\mathrm{N}}$ and all self-adjoint Laplacian realisations dominating $\mathcal{H}_{\mathrm{N}}$ in the sense of forms, including $\mathcal{H}_{\mathrm{F}}$. Since discreteness of the spectrum of self-adjoint realisations is equivalent to compactness of the embedding of their form domains in $L^{2}(\mathcal{G})$ this subsection can also be understood as discussing Sobolev embedding theorems. For (locally finite) infinite metric graphs with compact metric completion it has been known since [19, Theorem 2.2] that $H^{1}(\mathcal{G})$ embeds compactly into $C_{b}(\overline{\mathcal{G}})$, the space of bounded continuous functions over the metric completion of $\mathcal{G}$. We stress that, however, this does not necessarily imply a compact embedding into $L^{2}(\mathcal{G})$, unless $\overline{\mathcal{G}}$ has finite volume (for a counterexample, see the diagonal comb graph in 3.4.(2)).
In particular, the resolvent of $\mathcal{H}_{\mathrm{N}}$ and hence – by pointwise domination of the integral kernels – also of $\mathcal{H}_{\mathrm{F}}$ or any Markovian realisation – will be trace class operators, see [43, Theorem 5.1]. However, this finite volume criterion is insufficient to characterise discreteness of the spectrum: there are metric graphs of infinite volume on which $\mathcal{H}_{\mathrm{F}}$ and even $\mathcal{H}_{\mathrm{N}}$ have purely discrete spectrum: Non-trivial examples are given in [2, Theorem 1.3.5] and [62, Theorem 4.1].
@@
The subgraph $\mathcal{G}_{c}^{*}:=\mathcal{G}^{*}\setminus(\mathcal{G}_{c})^{\complement}$ can be identified with an Eulerian walk of finite length $p$ which allows to identify $f^{*}_{|\mathcal{G}_{c}^{*}}\in B^{*}$ with a function $\hat{f}\in H^{1}([0,p])$. On $\mathcal{G}_{c}^{*}$ we can then define
@@
(3.2) $\int_{\mathcal{G}_{c}^{*}}\lvert f^{*}(x+h)-f^{*}(x)\rvert^{2}\mathrm{d}x<\varepsilon^{2}.$
We cover $[0,p]$ with $N$ non-overlapping intervals of length $\delta$,
$0,p]=\bigcup_{j=1}^{N}\overline{I_{j}},\quad\text{where}\quad I_{j}\cap I_{k}=\emptyset,\quad\text{for $k\neq j$},\quad\text{and}\quad\lvert I_{j}\rvert=\delta\quad\text{for all $j\in\{1,\dots,N\}$}.$
@@
and note that the first summand in (3.3) is at most $\varepsilon^{2}$ by assumption. The second term in (3.3) can be considered as an integral along the Eulerian walk and estimated using Jensen’s inequality by
$\sum_{j=1}^{N}\int_{I_{j}}\left\lvert\frac{1}{\delta}\int_{I_{j}}\left(f^{*}(x)-f^{*}(y)\right)\mathrm{d}y\right\rvert^{2}\mathrm{d}x\leq\sum_{j=1}^{N}\int_{I_{j}}\frac{1}{\delta}\int_{I_{j}}\left\lvert f^{*}(y)-f^{*}(x)\right\rvert^{2}\mathrm{d}y\ \mathrm{d}x.$
We substitute h=yxh=y-x and emphasise that $h\in(-\delta,\delta)$. Thus, (3.2) implies
$\lVert f^{*}-Pf^{*}\rVert_{L^{2}(\mathcal{G^{*}})}^{2}$ $\leq\varepsilon^{2}+\sum_{j=1}^{N}\int_{I_{j}}\frac{1}{\delta}\int_{-\delta}^{\delta}\left\lvert f^{*}(x+h)-f^{*}(x)\right\rvert^{2}\mathrm{d}h\ \mathrm{d}x$
@@
Furthermore, $P$ is a bounded operator, and $B^{*}$ is bounded, thus $P(B^{*})$ itself is bounded. Finally, since $P$ is of finite rank, it is totally bounded. By [33, Lemma 1], we find that $B^{*}$ is totally bounded in $L^{2}(\mathcal{G}^{*})$, and thus the same is true of $B$ in $L^{2}(\mathcal{G})$.
@@
Next, we obtain a criterion on compactness of embeddings of Sobolev spaces on trees. It is a generalisation of [44, Lemma 8.1] where infinite trees without degree one vertices are considered. We also mention that the metric condition in Proposition 3.2 is already known to play a role for spectral properties of infinite quantum graphs with $\delta$-couplings, see [25, Theorem 3.5].
@@
On infinite trees without degree one vertices (“leaves”), the statement follows from [44, Lemma 8.1]
It remains to prove the statement in the presence of leaves.
Compactness of the embedding is equivalent to
@@ 3.2. Spectral phase transition on graphs of infinite volume
For $\alpha>0$, we define the “diagonal comb” metric graph $\mathcal{G}_{\alpha}$ by taking the interval $(0,1]$ (“horizontal shaft”), putting a vertex on every point $\frac{1}{n^{\alpha}}$, $n\in\mathbb{N}$, and attaching to it an edge (“tooth”) of length $\frac{1}{n^{\alpha}}$.
011
@@
Figure 3.1 contains an illustration in the case $\alpha=\frac{1}{2}$. The diagonal comb $\mathcal{G}_{\alpha}$ has, for any $\alpha\in(0,1]$, one topological end, diameter equal to 2 and volume $1+\sum\limits_{n=1}^{\infty}\frac{1}{n^{\alpha}}$. Note that the larger $\alpha$, the sparser the teeth become. In particular, the graph has finite volume if (and only if!) $\alpha>1$ in which case $\mathcal{H}_{\mathrm{N}}$ (or any Markovian realisation of $\mathcal{H}$) will have purely discrete spectrum by [19, Theorem 2.2].
@@
    If $\alpha\in(0,\frac{1}{2}]$, then there is an $L^{2}(\mathcal{G}_{\alpha})$-orthonormal sequence of $H^{1}(\mathcal{G}_{\alpha})$-functions with uniformly bounded $H^{1}(\mathcal{G}_{\alpha})$ norm. In particular, $\mathcal{H}_{\mathrm{N}}$ has non-empty essential spectrum.
@@
Before giving the proof of this result, a few observation are due. Note that the phase transition happens at $\alpha=1/2$, that is, among graphs of infinite volume (the transition from infinite to finite volume being at $\alpha=1$); see also Remark 3.6. Also, all $\mathcal{G}_{\alpha}$ are trees with only a finite number of edges of length larger than $\varepsilon$ for any $\varepsilon>0$ – the criterion for discreteness of the spectrum of the Friedrichs realisation in Proposition 3.2. Thus, Theorem 3.4 also demonstrates that this criterion for discreteness of the Friedrichs realisation on trees does not extend to the Kirchhoff realisation. Note that the operator described in (3) is not technically the Friedrichs realisation of the Laplacian in the sense of Definition 2.19 due to the presence of the additional Dirichlet conditions at the tips of the teeth.
@@
Denote by $\mathsf{e}_{k}$ the $k$-th tooth (counted from $1$ to 0, i.e. from right to left in the orientation of Figure 3.1), and note that all points in $\mathsf{e}_{k}$ are at most at distance $2k^{-\alpha}$ to 0; hence
@@
[…]
@@
Since, after passing to a subsequence, we can make the $\phi_{n}$ have mutually disjoint supports, up to renormalisation they yield an $L^{2}(\mathcal{G})$-orthonormal sequence of $H^{1}(\mathcal{G})$ functions with uniformly bounded $H^{1}(\mathcal{G})$-norm.
(3) This is a direct consequence of Proposition 3.2. ∎
@@
It is known that on any such domain, the spectrum of the Dirichlet Laplacian is purely discrete [59, 64]. Theorem 3.4.(3) or more generally the compactness criterion of Proposition 3.2 seem to be the corresponding analogues on metric graphs.
However, for the Neumann Laplacian on domains, the situation is different. Indeed, on connected domains of infinite volume, the essential spectrum of the Neumann Laplacian is always non-empty, see the appendix of [22], which also provides examples of horn-shaped domains of finite volume with non-empty spectrum.
Thus, Theorem 3.4.(1)–(2), which describes a phase transition from purely discrete to non-empty essential spectrum for the Neumann realisation among metric graphs of infinite volume, seems to present a new, metric-graph specific phenomenon which has no obvious equivalent among domains.
Remark 3.6.
Theorem 3.4.(2) asserts that the essential spectrum is non-empty but provides no further information on its structure. However, a closer look at the proof shows that, for $\alpha$ strictly below the critical threshold $\alpha=\frac{1}{2}$, the infimum of the spectrum is zero, since the test functions constructed in the proof will then have Rayleigh quotients converging to zero. It would be interesting to understand $\inf\sigma(\mathcal{H}_{\mathrm{N}})$ in the critical case $\alpha=\frac{1}{2}$.
@@
Example 3.8.
To show that Theorem 3.7 also generalises the result of Theorem 3.4.(1), we will apply it to the case of the comb graph $\mathcal{G}_{\alpha}$ (Definition 3.3), for given $\alpha>\frac{1}{2}$. Here, for large $k\in\mathbb{N}$ we can asymptotically bound $|A_{\frac{1}{k+1},\frac{1}{k}}(\gamma_{i})|$ from above as follows: the annulus will capture a segment of the horizontal shaft of length $\frac{1}{k}-\frac{1}{k+1}\sim\frac{1}{k^{2}}$, while it will also include a section of a number of teeth. To estimate this latter number, we need to know asymptotically how many integers $n$ are such that $\frac{1}{n^{\alpha}}\in[\frac{1}{k+1},\frac{1}{k}]$, that is, how many integers $n$ are between $k^{1/\alpha}$ and $(k+1)^{1/\alpha}$. Some simple analysis shows that this number behaves like $k^{\frac{1}{\alpha}-1}$ as $k\to\infty$. Estimating the length of each tooth from above by $\frac{1}{k}$, we can thus asymptotically bound the size of the annulus from above by
$|A_{\frac{1}{k+1},\frac{1}{k}}(\gamma_{i})|\lesssim k^{\frac{1}{\alpha}-1}\cdot\frac{1}{k}+\frac{1}{k^{2}}\sim k^{\frac{1}{\alpha}-2},$
which is of the form required by Theorem 3.7 if (and only if) $\alpha>\frac{1}{2}$.
@@ 3.3. Spectral approximation
Since $\mathcal{G}$ is connected, and the semigroups generated by both $\mathcal{H}_{\mathrm{F}}$ and $\mathcal{H}_{\mathrm{N}}$ are positive and irreducible, standard Perron–Frobenius theory, see e.g. [9, Proposition 4.12], implies that $\mu_{1}$ and $\lambda_{1}$ are necessarily simple. If $\mathcal{G}$ has finite volume, then $\mu_{1}=0$ with corresponding eigenfunctions being the constant functions, while $\lambda_{1}>0$, as we shall see below. In order to emphasise the dependence on $\mathcal{G}$, we will also write $\mu_{k}(\mathcal{G})$ and $\lambda_{k}(\mathcal{G})$. In the latter case, if there is any danger of ambiguity we will also include the Dirichlet vertex set, $\lambda_{k}(\mathcal{G})=\lambda_{k}(\mathcal{G},\mathfrak{V})$.
@@
On the other hand, given any $u\in H^{1}_{0}(\mathcal{G})$, by (3.9) and the definition of $H^{1}_{0}(\mathcal{G})$ as the closure in $H^{1}$ of $H^{1}_{c}(\mathcal{G})$, there exists a sequence of functions $u_{n}\in H^{1}_{0}(\mathcal{G}_{n})$ such that $u_{n}\to u$ in $H^{1}_{0}(\mathcal{G})$. Fix $k\geq 1$, write $\varphi_{j}:=\varphi_{j}(\mathcal{G})\in H^{1}_{0}(\mathcal{G})$ for the normalised eigenfunctions for $\lambda_{j}(\mathcal{G})$, $j=1,\ldots,k$, which will be our $u$ and for each $j$ choose $u_{j,n}\in H^{1}_{0}(\mathcal{G}_{n})$ such that $u_{j,n}\to\varphi_{j}$ in $H^{1}_{0}(\mathcal{G})$ (and hence also in $C(\overline{\mathcal{G}})$). Note that, by the dominated convergence theorem (using, e.g., $2|\varphi_{i}\varphi_{j}|\in L^{\infty}(\mathcal{G})$ as a dominating function, cf. [43, Lemma 3.2]), for all $i\neq j$,
@@
By induction on the trees $\mathcal{T}_{1},\ldots,\mathcal{T}_{j}$, it suffices to show that $\mu_{k}(\mathcal{G})\leq\mu_{k}(\mathcal{G}\setminus\mathcal{T}_{1})$ for all $k\in\mathbb{N}$. But this, in turn, follows from an argument completely analogous to the proof of [15, Theorem 3.10(1)] (with $r=1$): we first observe that clearly
@@
for all $k\in\mathbb{N}$, since the spectrum of the disjoint union of $\mathcal{T}_{1}$ and $\mathcal{G}\setminus\mathcal{T}_{1}$ is equal to the union of their spectra (counting multiplicities), and $\mu_{1}(\mathcal{T}_{1})=0\leq\mu_{k}(\mathcal{G}\setminus\mathcal{T}_{1})$ for all $k\in\mathbb{N}$. Next observe that $\mathcal{G}$ is formed from $\mathcal{T}_{1}\dot{\cup}(\mathcal{G}\setminus\mathcal{T}_{1})$ by gluing the two graphs together at the vertex $\mathsf{v}_{1}$; at the level of $H^{1}$-spaces, we have that $H^{1}(\mathcal{G})$ may be identified with the codimension one subspace of $H^{1}(\mathcal{T}_{1}\dot{\cup}(\mathcal{G}\setminus\mathcal{T}_{1}))$ consisting of those functions whose values satisfy an additional continuity condition at $\mathsf{v}_{1}$ (or rather its preimage vertices in $\mathcal{T}_{1}$ and $\mathcal{G}\setminus\mathcal{T}_{1}$ before gluing). The proof of [15, Theorem 3.4] may be repeated verbatim to give